# Crepe: A Mobile Screen Data Collector Using Graph Query

### v1

@@ 4.2. User Testing
    A 30-minute introductory interview to familiarize participants with the study and set up the Crepe app on their devices;
@@
Figure 4. Number of unique Instagram Story Ads collected by participants using the Crepe data collector. The top subplot displays data from Participants 1-3 over a 72-hour period, while the bottom subplot shows data from Participants 4-7 over a 24-hour period. The x-axis represents the time intervals in hours, and the y-axis indicates the number of Instagram Story Ads collected within each time interval.
@@
The figure consists of two vertically stacked subplots, both with multi-colored bar graphs. The top subplot, titled ”Participants 1-3 (72 hours),” displays data for three participants over a 72-hour period. Each participant’s data is represented by a different color: Participant 1 in blue, Participant 2 in purple, and Participant 3 in green. The bars indicate the number of unique Instagram Story Ads collected by each participant at various time intervals, with the time in hours along the x-axis and the number of ads on the y-axis. The bottom subplot, titled ”Participants 4-7 (24 hours),” presents data for four participants over a 24-hour period. The participants’ data are represented by different colors: Participant 4 in yellow, Participant 5 in orange, Participant 6 in dark blue, and Participant 7 in light green. Similar to the top subplot, the bars show the number of unique Instagram Story Ads collected by each participant at different time intervals, with the time in hours along the x-axis and the number of ads on the y-axis. Both subplots have clearly labeled x and y axes, and a legend is provided to identify the participants by their assigned colors. The figure allows for a visual comparison of the Instagram Story Ad data collected by different participants over their respective data collection periods.
4.3. Study Results
After initial data analysis, Crepe was able to collect 358 Instagram Story advertisements from all seven participants, with an average of 51 ads per participant (Figure 4). The collected data reveals the number of unique Instagram Story Ads encountered by each participant at regular time intervals throughout their respective data collection periods. The data allow us to observe the frequency and distribution of Instagram Story Ads over time, potentially providing insights into the advertising strategies employed by businesses on the platform. We shared the data we collected from each participant with them individually upon request.
@@ 6. Limitations and Future Work
We plan to open source Crepe to promote the ethical use of our novel approach in future academic research data collection. We also welcome contributions from the research community to Crepe to make it more ethical, robust, and adaptive to various data collection scenarios.
@@ Appendix A Appendix
A.1. Details of Our Best Query Selection Prompt
As described in Section 3.2.2, we used large language models to select the best Graph Query that most accurately matches the researcher’s data collection intention. Specifically, our prompt uses the following structure:
We defined this new graph query to locate relevant data on mobile UI structure. Now, we have generated the following potential queries:
@@ A.1. Details of Our Best Query Selection Prompt
Which one matches the following description the best?
[User description of the target data]
Please only return the index number, without any additional information, even periods or comma.
Example response format: 0
Example generated graph query candidates can be found in Figure 3 section 2. In the user description of the data to collect, in our user interfaces, we encourage the user to explain the target data’s relations with other content on screen, e.g., “Instagram Story advertisement title, the brand name above text Sponsored”.

### v2

@@ 4.4. Study 3: Evaluating System Impact for Users (Field Study)
    A 30-minute introductory interview to familiarize participants with the study and set up Crepe on their devices;
@@
Figure 7. Temporal patterns of Instagram story ads collected in Study 3 using Crepe. This scatter plot displays unique data from participants P1–P3, who chose a 72-hour study period based on their availability. Circle size corresponds to the number of entries collected at each time point (shown by the number within each circle). The distribution reveals how ad exposure varied across participants and throughout the three-day period, with some participants experiencing more consistent ad delivery while others encountered more sporadic patterns. All data was passively collected while participants engaged in their regular smartphone usage. The 24-hour study results for the rest of participants is shown in Figure 8.
@@
This image shows a scatter plot titled “Scatter Plot of Instagram Story Ads Collection (72-hour study)” with the subtitle “Dot size proportional to number of entries.” The plot displays data for three participants (P1, P2, and P3) over a 72-hour period. The y-axis represents the participants, labeled from P1 at the bottom to P3 at the top. The x-axis shows time in hours, ranging from 0 to approximately 70 hours. Each data point is represented by a colored circle where the size of the circle corresponds to the number of Instagram story ads collected, with larger circles indicating more entries. P1 (shown in blue circles) has data points scattered at hours 0 (8 entries), around hour 25 (1, 3, and 5 entries), and at hour 70 (1 and 2 entries). P2 (shown in purple circles) has more frequent data collection with clusters at hours 0-5 (with entries of 6, 2, 19, and 2), hours 10-15 (with 10 entries multiple times), hours 20-30 (with 4, 7, and 4 entries), hour 45 (with 10 entries), and hour 65 (with 2 entries twice). P3 (shown in green circles) has data points at hours 0-5 (15 entries), hours 10-15 (9 and 10 entries), hours 40-45 (2 entries twice), hours 55-60 (7, 1, and 1 entries), and hour 65 (7 entries). The legend in the top right identifies each participant with their corresponding color: P1 (blue), P2 (purple), and P3 (green).
Figure 8. The distribution of Instagram story ads collected in Study 3 using Crepe. This scatter plot represents unique data collected from participants P4-P7, who chose a 24-hour study period based on their availability. Each circle represents a collection event, with size proportional to the number of ads captured (indicated by the number inside each circle). Participants installed Crepe on their mobile phones and continued their normal daily activities while the system collected Instagram story ads in the background.
[…]
This image displays a scatter plot titled “Scatter Plot of Instagram Story Ads Collection (24-hour study)” with the subtitle “Dot size proportional to number of entries.” The plot tracks data collected from four participants (P4, P5, P6, and P7) over a 24-hour period. The horizontal axis represents time in hours (from 0 to 25), and the vertical axis shows the four participants. Each data point appears as a colored circle — olive for P4, orange-brown for P5, dark blue for P6, and light green for P7. The size of each circle corresponds to the number of Instagram story ads collected at that time point, with larger circles representing more entries. The specific number of entries is displayed within each circle. Participant P7 (light green circles) has the largest single collection of 35 entries at hour 0 and a smaller collection of 2 entries around hour 15. Participant P6 (dark blue circles) shows consistent activity throughout the 24-hour period, with multiple collection points at hours 0-2, 15-16, 18, 20, and 23. Participant P5 (orange-brown circles) shows activity primarily in the first 5 hours, with entries of 3-10 ads collected. Participant P4 (olive circles) shows a pattern of data collection at the beginning (hour 0), middle (hour 10), and end (hours 18-22) of the study period, with the largest collections of 24 and 28 entries occurring in the latter half of the study.
Study Results
After initial data analysis, Crepe was able to collect 358 Instagram Story advertisements from all seven participants, with an average of 51 ads per participant (Figure 7, Figure 8). The collected data reveals the number of unique Instagram Story Ads encountered by each participant at regular time intervals throughout their respective data collection periods. The data allow us to observe the frequency and distribution of Instagram Story Ads over time, potentially providing insights into the advertising strategies employed by businesses on the platform. We shared the data we collected from each participant with them individually upon request.
@@
4.5. Summary
Our three-part evaluation of Crepe examined its effectiveness in supporting researchers’ data collection needs (RQ1), accuracy in collecting target screen data (RQ2), and real-world performance impact (RQ3). Study 1 demonstrated that researchers could effectively use Crepe’s programming by demonstration approach, with all five participants successfully creating accurate data collectors for three different scenarios and rating the system’s usability favorably. Study 2 validated Crepe’s technical robustness through an in-lab evaluation, achieving high collection accuracy with an overall F1 score of 96.0% across Instagram, Uber, and Chrome applications, though some challenges emerged with complex UI screens containing over 500 elements. Finally, Study 3’s field deployment with seven participants showed that Crepe had minimal impact on device performance, with most participants reporting battery usage of 2% or less and unanimously indicating no noticeable interference with their regular phone usage. The system successfully collected 358 Instagram Story advertisements during the field study, though we identified opportunities for improving background process management to prevent occasional data collection interruptions. Together, these results demonstrate that Crepe effectively balances researcher needs, technical accuracy, and user experience in mobile screen data collection, while highlighting specific areas for future enhancement.
@@ 5.1. Collection Support Tools for De-centralized Data Access
The democratization of screen information access breaks the “data monopoly” that operating systems providers and app developers have long maintained (Jia-Jun Li et al., 2022). It shifts the data ownership back to end users, instead of leaving it only concentrated in the hands of tech platforms. By helping users see their own data, we can enable customized features that monitor and understand their behavior, supporting more informed device and app usage. This transparency allows individuals to make better decisions about their digital habits and identify patterns they might want to change.
By providing academic researchers access to screen data, they can better audit algorithms without depending on platform APIs that companies can (and often do) remove at will (Ribeiro et al., 2020; Sandvig et al., 2014; Calacci and Pentland, 2022; Kaplan and Sapiezynski, 2024). This independent access creates a more sustainable research ecosystem that isn’t vulnerable to corporate gatekeeping. Researchers can investigate questions about algorithmic bias, content recommendation systems, or user experience issues with greater freedom and thoroughness. This approach represents a fundamental shift in how we think about digital data ownership, treating screen data as something that belongs primarily to users rather than platforms. It encourages a more balanced relationship between tech companies, users, and the research community.
@@ 6. Limitations and Future Work
Taking inspiration from previous PBD studies (Lieberman et al., 2000; Cypher, 1991), it is also helpful to ask users demonstrate on multiple target data examples. We will continuously incorporate these features into Crepe and welcome the research community to collaboratively improve it in the future. We plan to open source Crepe to promote the ethical use of our novel approach in future academic research data collection. We also welcome contributions from the research community to Crepe to make it more ethical, robust, and adaptive to various data collection scenarios.
@@ Appendix A Appendix
A.1. Details of Our Graph Query Translation Prompt
As described in Section 3.2.2, we used large language models to translate the candidate graph queries to natural language. Specifically, our prompt uses the following structure:
We defined a Graph Query to locate target data on mobile UI structure. It describes the unique attributes of the data we are targeting. For example, the query
(conj (HAS\_CLASS\_NAME android.widget.FrameLayout) (RIGHT (conj (hasText 6) (HAS\_CLASS\_NAME android.widget.TextView) (HAS\_PACKAGE\_NAME com.ubercab)) ) (HAS\_PACKAGE\_NAME com.ubercab))
stands for: the information that is located to the right of a text “6”
Below I have a few queries, can you help me translate them to human-readable format like above?
@@ A.1. Details of Our Graph Query Translation Prompt
Leave out UI element names (TextView, FrameLayout) and do not make any reference to “view”. Users only care about the data and information contained in the view instead of the UI elements themselves. “With numeric index xx” should be translated into “the xx in the list” (first, second, etc.). Be as concise as possible. Make sure you return the translation in the order I presented the queries above, separated by new lines. Return nothing else.
A.2. Evaluation Study 1 Participant Demographics
Here we present the demographics of our evaluation Study 1 participants, as discussed in Section 4.2.
ID Age Gender Research Experience Collected Data Types Data Collection Tools Data Collection Projects
P1 23 Male 3 years IDE logs, eye tracking, semi-structured interview, survey IDE plugin, screen recorder, audio recorder, eye tracker 2
P2 30 Female 5 years app usage, surveys, interviews Google Analytics, Amazon Web Services, think-aloud, interviews, surveys, Qualtrics, MAXQDA About 6 (5–10)
P3 26 Male 4 UI interaction, think-aloud, interview Computer program, recorder 4
P4 32 Female 2 years Interviews Audio recording 2
P5 26 Female 3.5 years Gaze, image description Eye tracker 2
Table 4. The demographics of our evaluation Study 1 participants, as discussed in Section 4.2.
[…]
This table summarizes demographic and research background information for five participants in Evaluation Study 1. Each row represents a participant (P1 to P5), and columns include age (ranging from 23 to 32), gender (three female, two male), and years of research experience (from 2 to 5 years). The table also lists the types of data each participant has collected in past projects, such as IDE logs, eye tracking, interviews, surveys, and gaze data. Tools used for data collection include screen recorders, audio recorders, IDE plugins, eye trackers, and online survey platforms. The final column reports the number of data collection projects each participant has worked on, ranging from 2 to about 6.
A.3. Evaluation Study 2 Participant Demographics
Here we present the demographics of our evaluation Study 1 participants, as discussed in Section 4.3.
ID Age Gender Instagram Frequency Uber Frequency Chrome Mobile Frequency
PB1 21 Male Every week Once in a few weeks Everyday
PB2 25 Male Everyday Once in a few months Everyday
PB3 23 Male Every week Every week Everyday
PB4 30 Male Everyday Once in a few weeks Everyday
PB5 26 Female Everyday Once in a few weeks Everyday
Table 5. Demographics information and app usage frequency for our evaluation Study 2 participants in Section 4.3.
[…]
This table presents demographic and app usage frequency data for five participants in Evaluation Study 2. Each row corresponds to a participant (PB1 to PB5), with columns for age (from 21 to 30), gender (three male, two female), and self-reported frequency of using three mobile apps: Instagram, Uber, and Chrome. Instagram usage ranges from “Every week” to “Everyday,” Uber usage varies from “Every week” to “Once in a few months,” and all participants reported using Chrome on mobile “Everyday.” The table provides context on participants’ familiarity with the apps used during the study.